OpenStudy (anonymous):

A parking garage charges $5.50 to park for 4 hours and $7.75 to park for 7 hours. If the cost is a linear function of the number of hours parked, what is the cost to park for 3 hours?

OpenStudy (photon336):

\[y = m(x)+b \] you can find the rate of change or the slope, I.E. how much money is charged per hour. do you know how to do this?

OpenStudy (anonymous):

cost over time?

OpenStudy (photon336):

yep, in other words we can just build a function and get the slope

OpenStudy (photon336):

\[slope = \frac{ Change~~y }{ Change~~x }\] \[m = \frac{ 7.75-5.50 }{ 7-4 } = \frac{ 2.25 }{ 3 } = 0.75x\]

OpenStudy (photon336):

\[y = 0.75(x)\] so now we have our function. 0.75 cents per hour so now you can easily find the cost for three hours by inserting x = 3 into the equation.

OpenStudy (anonymous):

thank you

OpenStudy (photon336):

@rwally22 do you get what I did. does this make sense to you?

OpenStudy (anonymous):

yes, appreciate it